# SWAA: Sliding Window Attention Adaptation for Efficient Long-Context LLMs Without Pretraining

### v3

@@ 5.3 Performance–efficiency Trade-offs and Recommended Recipes
Although integrating more methods can typically achieve higher accuracy, it introduces more overhead, indicating that the efficiency of each recipe must also be evaluated. To assess the performance-efficiency trade-off of different SWAA configurations, we evaluate time-to-first-token (TTFT), time-per-output-token (TPOT), total throughput, and average running time per request. Concretely, we benchmark Qwen3-4B-Thinking on a single H100 GPU using vLLM’s bench_serve utility kwon2023efficientmemorymanagementlarge with random input data and 100 total requests. The prompt length and output length are set to 128k and 512 tokens, respectively, representing a typical long-context QA setting.
To visualize the performance-efficiency trade-off, Figure 2 plots each configuration’s accuracy on LongMemEval_24k wu2025longmemevalbenchmarkingchatassistants against its average running time, while detailed TTFT, TPOT, and throughput statistics for each configuration are provided in Appendix D. We draw a line between the full-attention point and the naive-SWA point as a baseline curve: configurations above this line offer a better accuracy-latency balance intuitively. For configurations with nearly identical time costs, we display only the one with the highest accuracy. Since Keep First $k$ has negligible impact on runtime (Appendix D), all plotted configurations fix $k=10$.
@@ 7 Limitations
We hypothesize that the ideal reasoning trajectory of the model adapted to SWA should be longer than the original model, to compensate for the information loss caused by SWA. That means, using the answers generated by the original model as fine-tuning data may not be the optimal training method. Rather, RL methods like GRPO shao2024deepseekmathpushinglimitsmathematical might further help the model adapted to SWA by learning a more adaptive reasoning trajectory. However, we did not experiment with them since they are too time-consuming and unstable.
@@ Appendix B Some Problems of Other Long-context Benchmarks
    LongBench bai2024longbenchbilingualmultitaskbenchmark is classic and widely used, but its average context length (most are under 16k) is relatively short for modern models, i.e., it is already too easy. And its data source is too old, leading to a risk of test data leakage. So we choose not to use it.
@@
    Rulerhsieh2024rulerwhatsrealcontext has controllable context length, but its tasks are almost all synthetic and most of them are needle-retrieval tasks, thus failing to reflect the model’s overall long-context capability in real-world scenarios.
@@
    LongBench-V2 bai2025longbenchv2deeperunderstanding is well-designed to necessitate deep understanding and reasoning over very long context. But it is too challenging for 4B-level models (e.g., Qwen3-4B-Thinking only gets 35% accuracy, which is too close to the random guessing baseline of 25%), making the improvement of different methods less distinguishable. Moreover, since it is in a multiple-choice question format, the results may not be sufficiently reliable because the model has a 25% chance of guessing the correct option.
    However, despite these drawbacks, they have been widely used in long-context model benchmarking. Thus we still elect to conduct our experiments on LongBench V2 bai2025longbenchv2deeperunderstanding and Ruler hsieh2024rulerwhatsrealcontext to verify the generalizability of our conclusions, as shown in Appendix C.
@@ Appendix F Results of LightTransfer
LightTransfer zhang2025lighttransferlongcontextllmsecretly represents a promising attempt at SWA adaptation on full-attention models without pretraining. It proposes a layer selection method for SWA adaptation that calculates a "lazy ratio," represented by the ratio of attention from tokens at the end of the sequence (from a calibration dataset) to recent tokens versus global tokens. Layers with a higher "lazy ratio" are selected to apply SWA, while the rest retain full attention. This method is intuitive and theoretically sound, but our experiments reveal some negative results.
Since the complete code of LightTransfer is not open-source, we reproduce this method using LongAlign bai2024longalignrecipelongcontext as the calibration dataset for lazy layer detection, where the number of last tokens is set to 64, and the recent token window is set to 1024. From our experimental results shown in Table 8, we find that:

### v4

@@ 5.3 Performance–efficiency Trade-offs and Recommended Recipes
Although integrating more methods can typically achieve higher accuracy, it introduces more overhead, indicating that the efficiency of each recipe must also be evaluated. To assess the performance-efficiency trade-off of different SWAA configurations, we evaluate time-to-first-token (TTFT), time-per-output-token (TPOT), total throughput, and average running time per request. Concretely, we benchmark Qwen3-4B-Thinking on a single H100 GPU using vLLM’s bench_serve utility Kwon et al. (2023) with random input data and 100 total requests. The prompt length and output length are set to 128k and 512 tokens, respectively, representing a typical long-context QA setting.
To visualize the performance-efficiency trade-off, Figure 2 plots each configuration’s accuracy on LongMemEval_24k Wu et al. (2024) against its average running time, while detailed TTFT, TPOT, and throughput statistics for each configuration are provided in Appendix D. We draw a line between the full-attention point and the naive-SWA point as a baseline curve: configurations above this line offer a better accuracy-latency balance intuitively. For configurations with nearly identical time costs, we display only the one with the highest accuracy. Since Keep First $k$ has negligible impact on runtime (Appendix D), all plotted configurations fix $k=10$.
@@ 7 Limitations
We hypothesize that the ideal reasoning trajectory of the model adapted to SWA should be longer than the original model, to compensate for the information loss caused by SWA. That means, using the answers generated by the original model as fine-tuning data may not be the optimal training method. Rather, RL methods like GRPO Shao et al. (2024) might further help the model adapted to SWA by learning a more adaptive reasoning trajectory. However, we did not experiment with them since they are too time-consuming and unstable.
@@ Appendix B Some Problems of Other Long-context Benchmarks
    LongBench Bai et al. (2023) is classic and widely used, but its average context length (most are under 16k) is relatively short for modern models, i.e., it is already too easy. And its data source is too old, leading to a risk of test data leakage. So we choose not to use it.
@@
    RulerHsieh et al. (2024) has controllable context length, but its tasks are almost all synthetic and most of them are needle-retrieval tasks, thus failing to reflect the model’s overall long-context capability in real-world scenarios.
@@
    LongBench-V2 Bai et al. (2024b) is well-designed to necessitate deep understanding and reasoning over very long context. But it is too challenging for 4B-level models (e.g., Qwen3-4B-Thinking only gets 35% accuracy, which is too close to the random guessing baseline of 25%), making the improvement of different methods less distinguishable. Moreover, since it is in a multiple-choice question format, the results may not be sufficiently reliable because the model has a 25% chance of guessing the correct option.
    However, despite these drawbacks, they have been widely used in long-context model benchmarking. Thus we still elect to conduct our experiments on LongBench V2 Bai et al. (2024b) and Ruler Hsieh et al. (2024) to verify the generalizability of our conclusions, as shown in Appendix C.
@@ Appendix F Results of LightTransfer
LightTransfer Zhang et al. (2024) represents a promising attempt at SWA adaptation on full-attention models without pretraining. It proposes a layer selection method for SWA adaptation that calculates a "lazy ratio," represented by the ratio of attention from tokens at the end of the sequence (from a calibration dataset) to recent tokens versus global tokens. Layers with a higher "lazy ratio" are selected to apply SWA, while the rest retain full attention. This method is intuitive and theoretically sound, but our experiments reveal some negative results.
Since the complete code of LightTransfer is not open-source, we reproduce this method using LongAlign Bai et al. (2024a) as the calibration dataset for lazy layer detection, where the number of last tokens is set to 64, and the recent token window is set to 1024. From our experimental results shown in Table 8, we find that: